OpenStudy (anonymous):

If sin a and cos a are the roots of px^2 + qx + r=0 then..which is correct p^2 - q^2 +2pr=0 p^2 + q^2 - 2pr=0

OpenStudy (anonymous):

now can you do it???

OpenStudy (anonymous):

i think i can plzz wait

OpenStudy (anonymous):

@THE_PROPHET i thinku have written it wrong

OpenStudy (anonymous):

kkkk....

OpenStudy (anonymous):

i get it sorry!!

OpenStudy (anonymous):

Can u correct it!

OpenStudy (anonymous):

\[(\sin \theta + \cos \theta)^{^{2}} = (\alpha + \beta)^{2}\]

OpenStudy (anonymous):

first substitute sina in place of x.....get an equation.... the substitute cos a ...get an equation......add both of them..... psin2a+qsina+x=0 pcos2a+qcosa+r=0

OpenStudy (anonymous):

\[\sin \theta X \cos \theta = r/p\]

OpenStudy (anonymous):

adding both of them... P+q(sina +cosa)+2r=0 but sina+cos a =-q/p(sum of roots) p+q(-q/p)+2r=0 p^2-q^2+2rp=o

OpenStudy (anonymous):

@THE_PROPHET i cant understand

OpenStudy (anonymous):

because sina and cosa ARE ROOTS OF THE EQUATION....they satisfy the equtaion.... so we can substitute sina and cosa in place of x to get the same value of 0....